OpenStudy (anonymous):

How do you solve 2/3k - (k + 1/4) = 1/12 (k + 4)

OpenStudy (anonymous):

k= -2 1/10

OpenStudy (anonymous):

That's not right

OpenStudy (anonymous):

its 1/12 instead of 1/2 sorry

OpenStudy (anonymous):

Sorry I was looking at a different problem yea it is

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

You should simplify it to 2/3k-k-1/4=1/12k+1/3

OpenStudy (anonymous):

Couldn't you eliminate the fractions

OpenStudy (anonymous):

You then subtract 2/3k from k so it would go to -1/3 k

OpenStudy (anonymous):

then you would change it into -4/12k so they are like factors and add -4/12 k to both sides so then you have -1/4=5/12k+1/3

OpenStudy (anonymous):

Is the answer -1/5

OpenStudy (anonymous):

.???

OpenStudy (anonymous):

no its -7/5